# Evolutionary Tracks and Spectral Properties of Quasi-stars and Their Correlation with Little Red Dots

### v1

@@ 2.3 Host Galaxy & Dust Corrections
Multiple explanations for the observed LRD spectra have been proposed. Leading theories are that these are compact, luminous objects (such as a quasi-star or black hole star) with additional components in the emission deriving from host galaxies or dust, although dust is presently disfavored (de Graaff, Anna et al., 2025; I. Labbe et al., 2024; R. P. Naidu et al., 2025; D. J. Setton et al., 2025). Observations are presently inconclusive with regard to the issue of the contribution to the flux from the stellar component of the host galaxies. Additionally, dust properties may vary between sources and with redshift, so we remain agnostic when exploring the impact of the stellar population of a host galaxy and dust to our models.
@@
In order to combine our quasi-star model with that of its host galaxy, we use spectral data from the Dawn JWST Archive (DJA, de Graaff, Anna et al., 2025; K. E. Heintz et al., 2024; F. Valentino et al., 2023). We create semi-synthetic models by simply combining our theoretical model with an observed star-forming galaxy at a redshift similar to the observed LRD. The hosts we have selected (UNCOVER-24996 and MoM-UDS-948311) primarily contribute to the UV spectrum where they dominate, and are found at similar sky positions and redshifts to their respective LRDs ( and ) (R. P. Naidu et al., 2024, 2025). In the quasi-star scenario, a host similar to those selected would serve as the ionizing source that prevents the fragmentation due to molecular hydrogen, thus enabling the formation of the quasi-star (B. Agarwal et al., 2013, 2016; M. C. Begelman et al., 2006; P. Natarajan et al., 2017; R. P. Naidu et al., 2025; J. A. Regan et al., 2017; C. Shang et al., 2010; E. Visbal & Z. Haiman, 2018).
@@
We implement dust corrections to our models using the simplified methods of S. Salim & D. Narayanan (2020) via
@@
where is the dust-free model (i.e. the quasi-star on its own) and is the dust curve itself. Although there are several components that go into calculating , we use a simple case based only on the UV-optical slope and a small UV-bump with strength . In this work, we use optical depth and bump strength . It should be noted that the effects of this bump strength are minimal and can be safely ignored if necessary, and that values this high have been ruled out for several LRDs (D. J. Setton et al., 2025).
3 Results
@@
Here we compare the predicted spectra of a fiducial M⊙ quasi-star model with JWST observations of two LRDs: UNCOVER-45924 and MoM-BH*-1 (I. Labbe et al., 2024; R. P. Naidu et al., 2025). We present the former as an example of a “typical” LRD and the latter as an example with an extreme Balmer break and only small contributions from a host galaxy. To facilitate this comparison, we normalize the flux at wavelengths corresponding to continuum features, and correct the wavelengths in the model and observations for the redshift in order to compare in the rest frame.
@@
Although this quasi-star model was not specifically fit to either LRD, it qualitatively reproduces the main continuum features seen in both. In order to make this comparison more quantitative, and despite the aforementioned limitations of our adopted atmospheric models which prevent detailed comparisons of individual spectral features, we compare the slopes of the UV and NIR-approach as well as the strength of the Balmer break. We find general agreement, with a few caveats listed below. Note that although the same quasi-star model is used for both comparisons, the different normalizations cause the model to have different values when compared to different sources.
@@
For each of our configurations (i.e. a quasi-star with and without a host, each with and without dust), we calculate the UV and NIR-approach slope parameters (where ) by performing linear regression in log-log space over the rest-frame wavelength ranges 0.18–0.28 m and 0.51–0.57 m, respectively. We then do the same for each LRD and compare the results. We find that the comparisons are in good agreement () except in two cases, both of which are in the UNCOVER comparison with dust: with a host in the UV region () and without a host in the NIR-approach (). This is concordant with other studies finding minimal evidence of dust attenuation in LRDs (D. J. Setton et al., 2025).
@@
We calculate the Balmer break strengths using the mean flux in the wavelength windows of Å and Å. For the UNCOVER source, our model yields a break strength of , closely matching the LRD’s value of . The value of the break in our model does not change significantly with the addition of a host galaxy, as the Balmer break window is dominated by the quasi-star flux. Although the far-UV portion of the standalone quasi-star is relatively weak, we obtain the characteristic LRD “V”-shape with the addition of a host galaxy. The inclusion of dust in the model slightly alters the fit by maintaining the NIR continuum. This modification has little effect, in agreement with recent claims that LRDs are mostly dust-free (C. M. Casey et al., 2025; R. P. Naidu et al., 2025; D. J. Setton et al., 2025; Xiao, Mengyuan et al., 2025).
@@
The other source, MoM-BH*-1, has one of the most pronounced Balmer breaks () of any LRD discovered (R. P. Naidu et al., 2025). While this is a more pronounced break than in our model (), as noted previously, this comparison may improve significantly with the adoption of more appropriate atmosphere models. Furthermore, this comparison is with a fiducial M⊙ quasi-star model not specifically fit to this source; a higher-mass model may reproduce these features more closely. Based on Fig. 1 we can estimate that a quasi-star model containing a more massive black hole of a few M⊙ will closely match the LRD’s bolometric luminosity. These topics will be the subjects of future investigations.
4 Conclusions & Discussion
@@
We have presented new MESA quasi-star evolutionary models generated with the MESA-QUEST framework and compared their synthetic spectra to JWST observations of LRDs. Our results demonstrate that late-stage quasi-stars, with black holes comprising 10% of their total mass and photospheric temperatures near 6000 K, naturally produce the defining continuum features of LRDs: large Balmer breaks, flat or red near-infrared slopes, and weak ultraviolet emission. These properties emerge self-consistently from the quasi-star’s radiation-supported envelope, and do not require an old stellar population or extreme dust reddening.
@@
The quasi-star phase of direct collapse formation of heavy seed black holes offers a plausible solution to the rapid emergence of supermassive black holes within the first few hundred million years of the Universe. A population of short-lived ( 20 Myr), extremely luminous quasi-stars could account for the high observed number density of LRDs, implying that early SMBHs may assemble by passing through this stage. Folding in lifetimes from our models along with predictions of progenitor cloud masses will allow us to predict if this is the case. Given the short lifetimes in our models, this can plausibly show that all SMBHs formed this way.

### v2

@@ II.3 Host Galaxy & Dust Corrections
Multiple explanations for the observed LRD spectra have been proposed. Leading theories are that these are compact, luminous objects (such as a quasi-star or black hole star) with additional components in the emission deriving from host galaxies or dust, although dust is presently disfavored (de Graaff, Anna et al., 2025; Labbe et al., 2024; Naidu et al., 2025; Setton et al., 2025). Observations are presently inconclusive with regard to the issue of the contribution to the flux from the stellar component of the host galaxies. Additionally, dust properties may vary between sources and with redshift, so we remain agnostic when exploring the impact of the stellar population of a host galaxy and dust to our models.
@@
In order to combine our quasi-star model with that of its host galaxy, we use spectral data from the Dawn JWST Archive (DJA, dja1; dja2; dja3). We create semi-synthetic models by simply combining our theoretical model with an observed star-forming galaxy at a redshift similar to the observed LRD. The hosts we have selected (JADES-27368, MoM-UDS-948311, and UNCOVER-24996) primarily contribute to the UV spectrum where they dominate, and are found at sky positions and redshifts similar to their respective LRDs ( and respectively) (D’Eugenio et al., 2025; Naidu et al., 2024, 2025). In the quasi-star scenario, a host similar to those selected would serve as the ionizing source that prevents fragmentation due to molecular hydrogen, thus allowing the formation of the quasi-star (Agarwal et al., 2013, 2016; Begelman et al., 2006; Natarajan et al., 2017; Naidu et al., 2025; Regan et al., 2017; Shang et al., 2010; Visbal and Haiman, 2018).
@@
We implement dust corrections to our models using the simplified methods of Salim and Narayanan (2020) via
@@
where is the dust-free model (i.e. the quasi-star on its own) and is the dust curve itself. Although there are several components that go into calculating , we use a simple case based only on the UV-optical slope and a small UV-bump with strength . In this work, we use optical depth and bump strength . It should be noted that the effects of this bump strength are minimal and can be safely ignored if necessary, and that values this high have been ruled out for several LRDs (Setton et al., 2025).
III Results
@@
Here we compare the predicted spectra of a fiducial M⊙ quasi-star model with JWST observations of three LRDs: “The Cliff”, MoM-BH*-1, and UNCOVER-45924 (de Graaff, Anna et al., 2025; Labbe et al., 2024; Naidu et al., 2025). We present the former two as examples with extreme Balmer breaks and only small contributions from a host galaxy, and the latter as an example of a more “typical” LRD with significant host contributions. To facilitate this comparison, we normalize the flux at wavelengths corresponding to continuum features, and correct the wavelengths in the model and observations for the redshift in order to compare in the rest frame.
@@
Although the quasi-star models were not specifically fit to any LRD, they qualitatively reproduce the main continuum features seen in all three presented here. In order to make this comparison more quantitative, and despite the aforementioned limitations of our adopted atmospheric models which prevent detailed comparisons of individual spectral features, we compare the slopes of the UV and NIR-approach as well as the strength of the Balmer break. We find general agreement, with a few caveats listed below.
@@
For each of our configurations (i.e. a quasi-star with and without a host, each with and without dust), we calculate the UV and NIR-approach slope parameters (where ) by performing linear regression in log-log space over the rest-frame wavelength ranges 0.18–0.28 m and 0.51–0.57 m, respectively. We then do the same for each LRD and compare the results. We find that the comparisons without dust are in good agreement, particularly with when a host is included ( in the NIR, ). All cases that include dust have significant discrepancies in the NIR, concordant with other studies finding minimal evidence of dust attenuation in LRDs (Setton et al., 2025).
@@
We calculate the Balmer break strengths of our models and the LRDs using a variation on the the revised index from Wang et al. (2024). We use the mean flux in the same blue wavelength window ( Å) and change the red window from the original Å to Å in order to avoid the CaII absorption features in our models discussed in Sec. II.2. Note that this leads to slightly different values for the Balmer break strengths than reported in each LRD’s original publication. Our model yields a break strength of , closely matching the UNCOVER LRD’s value of . The value of the break in our model does not change significantly with the addition of a host galaxy, as the Balmer break window is dominated by the quasi-star flux. Although the far-UV portion of the standalone quasi-star is relatively weak, we obtain the characteristic LRD “V”-shape with the addition of a host galaxy. The inclusion of dust in the model reduces their similarities to the observed LRDs in all cases, deviating significantly in the NIR in particular. This is in agreement with recent claims that LRDs are mostly dust-free (Casey et al., 2025; Naidu et al., 2025; Setton et al., 2025; Xiao, Mengyuan et al., 2025).
@@
The sources “The Cliff” and MoM-BH*-1 have some of the most pronounced Balmer breaks ( and respectively) of any LRD discovered (Naidu et al., 2025). While this is a more pronounced break than in our model (), as noted previously, this comparison may improve significantly with the adoption of more appropriate atmosphere models. Furthermore, this comparison is with a fiducial M⊙ quasi-star model not specifically fit to this source; a higher-mass model may reproduce these features more closely. Based on Fig. 1 we can estimate that a quasi-star model containing a more massive black hole of a few M⊙ will closely match the LRD’s bolometric luminosity. These topics will be the subjects of future investigations.
IV Conclusions & Discussion
@@
We have presented new MESA quasi-star evolutionary models generated with the MESA-QUEST framework and compared their synthetic spectra to JWST observations of LRDs. Our results demonstrate that late-stage quasi-stars, with black holes comprising 10% of their total mass and photospheric temperatures near K, naturally produce the defining continuum features of LRDs: large Balmer breaks, flat or red near-infrared slopes, and weak ultraviolet emission. These properties emerge self-consistently from the quasi-star’s radiation-supported envelope, and do not require an old stellar population or extreme dust reddening. While some of these features may also appear in supermassive stars, they are expected to have both order of magnitude lower luminosities and up to several thousand K greater temperatures than those found in late-stage quasi-stars of equivalent masses, as well as extremely short lifetimes in which they form and collapse within years (Hosokawa et al., 2013; Nandal and Loeb, 2025).
@@
The quasi-star phase of direct collapse formation of heavy seed black holes offers a plausible solution to the rapid emergence of supermassive black holes within the first few hundred million years of the Universe. With lifetimes of Myr, quasi-stars would exist for roughly of the age of the universe at redshift . Combining this duty cycle with the observed LRD comoving density of cMpc-3 at similar redshifts (Greene et al., 2024), we estimate that the total number density of objects that undergo an LRD phase is cMpc-3. If quasi-stars have a median mass of M⊙ and subsequently accrete an additional factor of during their post-quasi-star evolution, the cumulative black hole mass density contributed by this population reaches M⊙cMpc-3. Optimistically, this value is only about a factor of two below the observed black hole mass density of M⊙cMpc-3 (Yu and Tremaine, 2002), suggesting that a population of short-lived, extremely luminous quasi-stars could account for a significant fraction of present-day supermassive black holes. Folding in lifetimes from our models along with predictions of progenitor cloud masses will allow us to predict if this is the case. Given the short lifetimes in our models, this can plausibly show that a significant fraction, or even all SMBHs formed this way.